OpenStudy (anonymous):

What is the discontinuity and zero of the function f(x) = the quantity of 3 x squared plus x minus 4, all over x minus 1?

OpenStudy (anonymous):

Discontinuity at (-1, 1), zero at (four thirds, 0) Discontinuity at (-1, 1), zero at (negative four thirds , 0) Discontinuity at (1, 7), zero at (four thirds , 0) Discontinuity at (1, 7), zero at (negtive four thirds, 0)

hartnn (hartnn):

\(\Large f(x) = \dfrac{3x^2+x-4}{x-1}\)

hartnn (hartnn):

to get the zeros, put f(x) = 0

hartnn (hartnn):

so, numerator = 0, \(3x^2+x-4 = 0\) xan you solve this quadratic ?

hartnn (hartnn):

*can

OpenStudy (anonymous):

let me see tell me if its right

OpenStudy (anonymous):

I have so far 9 + x - 4=0

hartnn (hartnn):

9 ? you've solve quadratics before ? using factor method

OpenStudy (anonymous):

whats next

OpenStudy (anonymous):

Don't I do 3 times 3?

hartnn (hartnn):

but thats not correct... \(3x^2 +x-4 = 0 \\ 3x^2 -3x +4x-4 = 0 \\ 3x (x-1)+4(x-1)=0\)

hartnn (hartnn):

ever done such thing ^^ ?

OpenStudy (anonymous):

ohh no lol

hartnn (hartnn):

\(3x^2 +x-4 = 0 \\ 3x^2 -3x +4x-4 = 0 \\ 3x (x-1)+4(x-1)=0 \\ (3x+4)(x-1) = 0\)

hartnn (hartnn):

thats called factoring the quadratic!

hartnn (hartnn):

see if you get the steps, ask if any doubts..

OpenStudy (anonymous):

ok I will

OpenStudy (anonymous):

whered the one come from in 3x(x - 1)

hartnn (hartnn):

\(3x^2 = 3\times x \times x \\ -3x = -3 \times x\) so 3 and 'x' are common, factoring them out! \(3\times x (x -1) = 3x (x-1)\)

hartnn (hartnn):

got it ?

hartnn (hartnn):

going forward, \((3x+4)(x-1) =0 \\ 3x+4 = 0 \quad or \quad x-1= 0\) but x-1 cannot be = 0 , since denominator cannot be = 0 so only 3x+4 = 0 what do you get x from here ?

hartnn (hartnn):

the function will be discontinuous at points where denominator = 0 so, for discontinuity solve x-1= 0, x =... ?